OpenStudy (anonymous):

an isosceles triangle has two angles that measure 75. what is the measure of the third angle of the triangle?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Hi.

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Okay so what is an Isosceles triangle?

OpenStudy (anonymous):

help me plz

OpenStudy (anonymous):

I'm will help.

OpenStudy (anonymous):

Isosceles Triangle: Two Sides Are Of Equal Length.

OpenStudy (anonymous):

Okay good. Give me a second so I can work the problem out, and I will explain how to do it.

OpenStudy (anonymous):

i dont know your need to get that

OpenStudy (anonymous):

180 - (75+75) = 180 - 150 = 30

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The third angle of the isosceles triangle is 30°.

OpenStudy (anonymous):

are you a boy or a girl

OpenStudy (anonymous):

The sum of the angles of a triangle is equal to 180°. Do you get it?

OpenStudy (anonymous):

Girl.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

byee

OpenStudy (anonymous):

Okay then.